# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v5

@@ 7.2 Spatial coherence
The inter-pixel dependency of segmentation masks often introduces additional complexity in proper uncertainty quantification, questioning the commonly repurposed techniques from conventional image-level classification. In this section, we further dive into import details and challenges arising due to the added dimension(s) in both 2D and 3D data.
@@
Quantifying the epistemic uncertainty requires subtracting the expected entropy from the total entropy of the predictive distribution (see equations (3) and (39)). In such cases, the entropy is usually estimated via mean or summation over all pixel-level uncertainty values Mukhoti & Gal (2018); Camarasa et al. (2021); Bhat et al. (2022b); Mukhoti et al. (2020); Ma et al. (2024b); Shen et al. (2021); Li & Alstrøm (2020); Dechesne et al. (2021); Huang et al. (2018); Nair et al. (2020); Zepf et al. (2023b). However, this asserts the assumption of a factorized categorical distribution, i.e. uncorrelated pixels, which is a strong assumption that ignores the core rationale for the field’s existence. Since this could explain failure cases of entropy-based uncertainty evaluation methods, it is important to shift towards contemporary generative models that enable exact likelihoods, improving the theoretical soundness of the evaluated entropies. Moreover, elements such as object class dependency, size, shape, proximity, contextual information etc. can all influence the decision-making of the aggregation method. Notably, Kahl et al. (2024) argue that pixel aggregation should be considered as a separate, distinct modeling choice in probabilistic segmentation. The authors demonstrate that naive summation-based aggregation techniques (Wang et al., 2019a; Jungo et al., 2020; Zhang et al., 2022a; Czolbe et al., 2021; Camarasa et al., 2021) can cause the foreground object size to correlate with the uncertainty score. Some empirical studies have explored aggregation (Camarasa et al., 2021; Roy et al., 2018b; Jungo et al., 2020; Kasarla et al., 2019; Xie et al., 2022; Wu et al., 2021). However, we advocate for deeper investigation into well-informed image-level aggregation strategies and their implications for uncertainty quantification.
@@
In clinical settings, data is often acquired volumetrically but sliced into patches due to memory limits. With improved compute, 3D segmentation models have gained traction for preserving spatial continuity and reducing inter-slice inconsistencies. As we have seen the implications of the additional dimension in 2D uncertainty, we similarly argue that 3D segmentation introduces or exacerbates previously mentioned issues related to entropy estimation and aggregation of pixels. Probabilistic segmentation has been employed on 3D data but often are straightforward extensions of patch-wise segmentation architectures (Viviers et al., 2023b; Chotzoglou & Kainz, 2019; Long et al., 2021b; Saha et al., 2021a; 2020; Hoebel et al., 2019). Limited diversity of uncertainty-based sample selection in Active Learning requires more focus, since 2D slices of 3D volumes are often highly redundant Shen et al. (2021); Burmeister et al. (2022). Furthermore, given the high computational cost of many architectures, we emphasize the need for efficiency with higher-dimensional input tensors. Furthermore, methodologies need to be developed in order to appropriately compare 2D- to 3D models. For example, the GED cannot be estimated volumetrically with 2D models. Hence, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to phase out patch-based approaches for volumetric data. Fortunately, research dedicated to probabilistic volumetric segmentation is prevalent and its successes underline the potential for this research direction Liu et al. (2025); Burmeister et al. (2022); Viviers et al. (2023b).
@@ 7.3 Standardization
The widespread success of contemporary machine learning models can be largely attributed to open-source initiatives and rigorous benchmarking practices. The field of probabilistic segmentation falls short on this. A key factor contributing to these diverging standards is that the literature is predominantly application-driven, uncertainty often considered merely an auxiliary tool, leading to a fragmented landscape where studies rely on varied datasets and methodologies without a shared evaluation protocol. This problem has been partly addressed by contemporary research attempting to standardize the field through a unified framework after identifying key pitfalls (Kahl et al., 2024). These practices are included in our recommended practices, as discussed in Section 6 per application.
Despite the aforementioned points, efforts to quantify observer variability show some implicit consensus. Notably, the recommendations of Kahl et al. (2024) diverge from these practices, suggesting a more rigorous examination of this application. For instance, for the widely used LIDC-IDRI dataset (Armato, 2011), we recommend following the preprocessing protocol of Baumgartner et al. (2019), which is supported by publicly available code and has been adopted in several subsequent studies. Similarly, the Cityscapes dataset (Cordts et al., 2016a), comprising street-view images from German cities, has been used to introduce artificial class-level label ambiguity (Kassapis et al., 2021; Kohl et al., 2019b; Gao et al., 2022). We advise using the label-switching parameters proposed by Kohl et al. (2019b), whose preprocessing scripts are also available. Other multi-annotated datasets, such as those from the QUBIQ Challenge (noa, ), include MRI and CT scans of various organs (Valiuddin et al., 2024b; Ji et al., 2021), although they follow less standardized preprocessing. The RIGA dataset for retinal fundus images (Almazroa et al., 2018) can be used as in Wu et al. (2022).
The evaluation metrics used across studies also diverge substantially. While GED and Hungarian Matching are commonly applied to assess model performance, key details—such as the choice of distance kernel, the number of predictive samples, and the handling of correct empty segmentations (which cause NaNs) vary widely. This inconsistency is particularly evident in the work of Zbinden et al. (2023), who attempt to identify common ground across benchmarks but find over 10 (!) different evaluation methods. Additionally, many reported results are not taken from original papers but from follow-up studies using them as baselines. See Table 6 for a comparison of benchmarks on the LIDC-IDRI dataset. We recommend using 16 predictive samples as the standard for evaluating the LIDC-IDRI (v2, see experiments section of Zbinden et al. (2023)) dataset. Additionally, correctly predicted empty segmentations should be assigned full score. Notably, the literature often overemphasizes metric improvement, which can incentivize models to simply replicate ground-truth masks rather than capture the underlying distribution. This risks undermining the original goal of predicting plausible, unseen segmentations, a textbook case of Goodhart’s law: “When a measure becomes a target, it ceases to be a good measure” and is further compounded by known biases in commonly used metrics (Kohl et al., 2019a; Zepf et al., 2023a; Valiuddin et al., 2024b; a). Hence, practitioners should remain cautious and complement quantitative evaluations with qualitative assessments from domain experts.
@@
PhiSeg (Baumgartner et al., 2019) $0.262\pm 0.000^{*}$ $0.586\pm 0.000^{*}$
SSN (Monteiro et al., 2020) $0.259\pm 0.000^{*}$ $0.558\pm 0.000^{*}$
@@
JProb. U-Net (Zhang et al., 2022d) - - $0.262\pm 0.000^{\phantom{*}}$ $0.585\pm 0.000^{\phantom{*}}$
PixelSeg (Zhang et al., 2022c) $0.243\pm 0.010^{\phantom{*}}$ $0.614\pm 0.000^{\phantom{*}}$ $0.260\pm 0.000^{\phantom{*}}$ $0.587\pm 0.010^{\phantom{*}}$
MoSE (Gao et al., 2022) $0.218\pm 0.001^{\phantom{*}}$ $0.624\pm 0.004^{\phantom{*}}$
AB (Chen et al., 2022) $0.213\pm 0.001^{*}$ $0.614\pm 0.001^{*}$
CIMD (Rahman et al., 2023) $0.234\pm 0.005^{*}$ $0.587\pm 0.001^{*}$
@@ 7.4 Future of Active Learning
Among the applications considered, uncertainty quantification emerges as a promising approach to Active Learning, showing true potential of potentially reducing annotation costs while preserving or even improving model generalization. Active Learning directly addresses the labeling bottleneck through uncertainty-driven selection, aligning with model introspection and fostering human-AI collaboration in specialized domains. Furthermore, when combined with federated learning, this field can accelerate privacy-centric collaboration by enabling efficient, human-in-the-loop improvement of safety-critical models across decentralized data silos. However, key limitations remain and require further research.
@@
While deep learning-based segmentation typically involves high-dimensional data and large datasets, active learning requires a well-generalized model trained on limited data. This mismatch renders performance highly sensitive to budget constraints, model architecture, hyperparameters, and regularization Mittal et al. (2019); Kirsch (2024); Munjal et al. (2022); Kirsch et al. (2019); Gaillochet et al. (2023). Furthermore, random sampling, particularly when combined with MC Dropout, often proves to be a surprisingly strong baseline that is difficult to outperform Kahl et al. (2024); Mittal et al. (2019); Lüth et al. (2023); Burmeister et al. (2022); Siddiqui et al. (2020); Kim et al. (2021); Sinha et al. (2019); Nath et al. (2020), especially with highly imbalanced data Ma et al. (2024a). This observation can partly be attributed to the fact that MC Dropout likely fails to capture true Bayesian uncertainty (see Section 7.7), inviting exploration of other techniques.
@@
Besides uncertainty, the diversity of selected samples is equally critical. For instance, a set of highly uncertain samples may share similar semantic content, leading the model to overfit a narrow and homogeneous region of the data distribution during early training. Hence, sample selection should be based on both uncertainty and diversity (Ozdemir et al., 2021; Wu et al., 2021; Jensen et al., 2019; Burmeister et al., 2022). Ozdemir et al. (2021) leverage the latent space of a VAE to select representative samples, whereas other studies use feature vectors from prediction models (Wu et al., 2021; Burmeister et al., 2022). Owing to their significantly greater representational capacity, modern generative models, as discussed in Section 7.6, merit experimental investigation together with segmentation specific conditioning. Furthermore, the sample diversity offered by such models can be effectively combined with adversarial training (Mahapatra et al., 2018; Sinha et al., 2019).
@@
As discussed in Section 4.1, using the predictive entropy of SoftMax-based uncertainty requires proper calibration. Unfortunately, this is rarely verified in practice Kasarla et al. (2019); García Rodríguez et al. (2020); Xie et al. (2022); Wu et al. (2021); Burmeister et al. (2022); Gaillochet et al. (2023). Furhtermore, most works rely on heuristics to aggregate the pixel-level information such as closeness to edges, boundaries or the use of super-pixels Gorriz et al. (2017); Kasarla et al. (2019); Ma et al. (2024b); Gaillochet et al. (2023); Wu et al. (2021); Kasarla et al. (2019); Bengar et al. (2021). Moreover, the mutual information is often quantified by plain summation (i.e. assuming pixel independence) Shen et al. (2021); Ma et al. (2024b). We hypothesize that these ad-hoc aggregation techniques, as mentioned in Section 7.2, are a key limiting factor to uncertainty-based Active Learning.
@@
Finally, the “cold start” problem Nath et al. (2022); Houlsby et al. (2014); Yuan et al. (2020a), the notion that uncertainty methods work poor on the initial samples, is rarely addressed. Most approaches assume a well-trained model is already available, yet methods like MC Dropout require warm-up epochs to yield meaningful uncertainty. Addressing this gap, post-hoc techniques such as that of Sourati et al. (2018), which enable uncertainty estimation from pretrained deterministic networks, represent a valuable and underexplored direction for early-stage sample selection.

### v8

@@ 7.2 Spatial coherence
The inter-pixel dependency of segmentation masks often introduces additional complexity in proper uncertainty quantification, questioning the commonly repurposed techniques from conventional image-level classification. In this section, we further dive into important details and challenges arising due to the added dimension(s) in both 2D and 3D data.
@@
The mutual information is often utilized to express the reducible, epistemic uncertainty. This requires subtracting the expected entropy from the total entropy of the predictive distribution (see equations (3) and (35)). In such cases, the entropy is usually estimated via mean or summation over all pixel-level uncertainty values (Mukhoti and Gal, 2018; Camarasa et al., 2021; Bhat et al., 2022b; Mukhoti et al., 2020; Ma et al., 2024b; Shen et al., 2021; Li and Alstrøm, 2020; Dechesne et al., 2021; Huang et al., 2018; Nair et al., 2020; Zepf et al., 2023b). However, this asserts the assumption of a factorized categorical distribution, i.e. uncorrelated pixels, which is a strong assumption that ignores the core rationale for the field’s existence. Hence, spatial coherence is a fundamental consideration and is frequently neglected in the uncertainty estimation frameworks used for Active Learning.
The problem with factorized entropy can be demonstrated with a simple example: a binary segmentation of a circle on a $\times32$ grid. If we assume that the ’latent’ uncertainty is purely translational, Figure 13 visualizes the example distributions of possible segmentation maps. The true entropy of the generating source scales with the number of possible states $N$ as $\log_{2}(N)$. However, using a factorized approach severely overestimates the latent, true uncertainty. This behavior is a direct consequence of the subadditivity property of entropy, which states that
$H[X_{1},X_{2},...,X_{D}]\leq\sum^{D}_{i}H[X_{i}].$ (43)
(a) 22 states, true entropy $log_{2}(2)=1$ bit, factorized entropy $p_{i}\sum\log_{2}p_{i}\approx 112$ bits
(b) 55 states, true entropy $log_{2}(2)\approx 2.32$ bits, factorized entropy $p_{i}\sum\log_{2}p_{i}\approx 114$ bits
(c) 88 states, true entropy $log_{2}(8)=3$ bits, factorized entropy $p_{i}\sum\log_{2}p_{i}\approx 190$ bits
Figure 13: This figure illustrates why factorized entropy is a poor proxy for the true uncertainty in structured images. The common method to estimate this entropy is by averaging multiple sampled binary segmentation masks, where the resulting pixel intensity indicates the average confidence value. The shapes shown possess only translational uncertainty. The factorized model fails to capture the object’s spatial coherence and leads to a significant overestimation of the true entropy.
Importantly, this fallacy could explain failure cases of entropy-based uncertainty evaluation methods in tasks such as Active Learning where sample informativeness/diversity is crucial. Elements such as object class dependency, size, shape, proximity, contextual information etc. can all influence the decision-making of the aggregation method. Notably, Kahl et al. (2024) argue that pixel aggregation should be considered as a separate, distinct modeling choice in probabilistic segmentation. The authors demonstrate that naive summation-based aggregation techniques (Wang et al., 2019a; Jungo et al., 2020; Zhang et al., 2022a; Czolbe et al., 2021; Camarasa et al., 2021) can cause the foreground object size to correlate with the uncertainty score. Some empirical studies have explored other types of aggregation (Camarasa et al., 2021; Roy et al., 2018b; Jungo et al., 2020; Kasarla et al., 2019; Xie et al., 2022; Wu et al., 2021). However, we advocate for deeper investigation into well-informed image-level aggregation strategies and their implications for uncertainty quantification.
@@
Processing volumetric clinical data knows two dominant strategies. The first, 2D slice processing, analyzes the volume as a stack of individual images, preserving full context within each slice but severing spatial continuity along the third dimension. The second, 3D patch processing, divides the volume into smaller ones, which maintains local 3D information but loses the global anatomical structure. As compute power increases, end-to-end 3D segmentation models are gaining traction because they can analyze increasingly more volume, preserving spatial continuity and reducing the inconsistencies inherent in these slice- or patch-based methods (Liu et al., 2025; Burmeister et al., 2022; Viviers et al., 2023c; Monteiro et al., 2020). As we have seen the implications of the additional dimension in 2D uncertainty, we similarly argue that 3D segmentation introduces or exacerbates previously mentioned issues related to entropy estimation and aggregation of pixels. Probabilistic segmentation has been employed on 3D data but often are straightforward extensions of patch-wise segmentation architectures (Viviers et al., 2023c; Chotzoglou and Kainz, 2019; Long et al., 2021a; Saha et al., 2021a; 2020; Hoebel et al., 2019). Limited diversity of uncertainty-based sample selection in Active Learning requires more focus, since 2D slices of 3D volumes are often highly redundant (Shen et al., 2021; Burmeister et al., 2022). Furthermore, given the high computational cost of many architectures, we emphasize the need for efficiency with higher-dimensional input tensors. Moreover, methodologies need to be developed in order to appropriately compare 2D to 3D models. For example, a valid volumetric GED cannot be estimated with 2D models. Hence, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to phase out patch-based approaches for volumetric data.
@@ 7.3 Standardization
The widespread success of contemporary machine learning models can be largely attributed to open-source initiatives and rigorous benchmarking practices. The field of probabilistic segmentation falls short on this. A key factor contributing to these diverging standards is that the literature is predominantly application-driven, uncertainty often considered merely an auxiliary tool, leading to a fragmented landscape where studies rely on varied datasets and methodologies without a shared evaluation protocol. This problem has been partly addressed by contemporary research attempting to standardize the field through a unified framework after identifying key pitfalls (Kahl et al., 2024). Our manuscript further contributes by introducing coherence and a unifying synthesis across the distinct disciplines of applications, tasks, and methodological developments, with these concepts being further detailed in our recommendations in Section 8.
Despite the aforementioned points, efforts to quantify observer variability show some implicit consensus. Notably, some of the recommendations of Kahl et al. (2024) diverge from these practices, such as how to employ the widely used LIDC-IDRI dataset (Armato, 2011) or the Cityscapes Cordts et al. (2016a) dataset, suggesting a more rigorous examination of this application. Furthermore, the evaluation metrics used across various studies also diverge substantially. While GED and Hungarian Matching are commonly applied to assess model performance, key details, such as the choice of distance kernel, the number of predictive samples, and the handling of correct empty segmentations (which cause NaNs), vary widely. This inconsistency is particularly evident in the work of Zbinden et al. (2023), who attempt to identify common ground across benchmarks but find over 10 (!) different evaluation methods. Additionally, many reported results are not taken from original papers but from follow-up studies using them as baselines. See Table 6 for a comparison of benchmarks on the LIDC-IDRI dataset.
@@
PhiSeg (Baumgartner et al., 2019) $0.262\pm 0.000^{*}$ $0.586\pm 0.000^{*}$ - -
SSN (Monteiro et al., 2020) $0.259\pm 0.000^{*}$ $0.558\pm 0.000^{*}$ - -
@@
JProb. U-Net (Zhang et al., 2022b) - - $0.262\pm 0.000^{\phantom{*}}$ $0.585\pm 0.000^{\phantom{*}}$
PixelSeg (Zhang et al., 2022d) $0.243\pm 0.010^{\phantom{*}}$ $0.614\pm 0.000^{\phantom{*}}$ $0.260\pm 0.000^{\phantom{*}}$ $0.587\pm 0.010^{\phantom{*}}$
MoSE (Gao et al., 2022) $0.218\pm 0.001^{\phantom{*}}$ $0.624\pm 0.004^{\phantom{*}}$ - -
AB (Chen et al., 2022) $0.213\pm 0.001^{*}$ $0.614\pm 0.001^{*}$ - -
CIMD (Rahman et al., 2023) $0.234\pm 0.005^{*}$ $0.587\pm 0.001^{*}$ - -
@@ 7.4 Future of Active Learning
Among the applications considered, uncertainty quantification emerges as a promising approach to Active Learning, showing true potential of potentially reducing annotation costs while preserving or even improving model generalization. Active Learning directly addresses the labeling bottleneck through uncertainty-driven selection, aligning with model introspection and fostering human-AI collaboration in specialized domains. Furthermore, when combined with federated learning, this field can accelerate privacy-centric collaboration by enabling efficient, human-in-the-loop improvement of safety-critical models across decentralized data silos. Similarly, these principles can also translate tp Reinforcement Learning, where uncertainty guides an agent’s exploration policy to accelerate learning in complex environments. However, key limitations remain and require further research.
@@
Active Learning requires a well-generalized model trained on limited data, rendering performance highly sensitive to budget constraints, model architecture, hyperparameters, and regularization (Mittal et al., 2019; Kirsch, 2024; Munjal et al., 2022; Kirsch et al., 2019; Gaillochet et al., 2023). Furthermore, random sampling, particularly when combined with MC Dropout, often proves to be a surprisingly strong baseline that is difficult to outperform (Kahl et al., 2024; Mittal et al., 2019; Lüth et al., 2023; Burmeister et al., 2022; Siddiqui et al., 2020; Kim et al., 2021; Sinha et al., 2019; Nath et al., 2020), especially with highly imbalanced data (Ma et al., 2024a). This observation can partly be attributed to the fact that MC Dropout likely fails to capture true Bayesian uncertainty (see Section 7.5), inviting exploration of other techniques.
@@
A set of highly uncertain samples may share similar semantic content, leading the model to overfit a narrow and homogeneous region of the data distribution during early training. Hence, sample selection should be based on both uncertainty and diversity (Ozdemir et al., 2021; Wu et al., 2021; Jensen et al., 2019; Burmeister et al., 2022). Ozdemir et al. (2021) leverage the latent space of a VAE to select representative samples, whereas other studies use feature vectors from prediction models (Wu et al., 2021; Burmeister et al., 2022). Furthermore, the sample diversity offered by such models can be effectively combined with adversarial training (Mahapatra et al., 2018; Sinha et al., 2019).
@@
As discussed in Section 4.1, using the predictive entropy of SoftMax-based uncertainty requires proper calibration. Unfortunately, this is rarely verified in practice (Kasarla et al., 2019; García Rodríguez et al., 2020; Xie et al., 2022; Wu et al., 2021; Burmeister et al., 2022; Gaillochet et al., 2023). Furhtermore, most works rely on heuristics to aggregate the pixel-level information such as closeness to edges, boundaries or the use of super-pixels (Gorriz et al., 2017; Kasarla et al., 2019; Ma et al., 2024b; Gaillochet et al., 2023; Wu et al., 2021; Kasarla et al., 2019; Bengar et al., 2021). Moreover, the mutual information is often quantified by plain summation (i.e. assuming pixel independence) (Shen et al., 2021; Ma et al., 2024b). We hypothesize that these ad-hoc aggregation techniques, as mentioned in Section 7.2, are a key limiting factor to uncertainty-based Active Learning.
@@
Finally, the “cold start” problem (Nath et al., 2022; Houlsby et al., 2014; Yuan et al., 2020a), the notion that uncertainty methods work poor on the initial samples, is rarely addressed. Most approaches assume a well-trained model is already available, yet methods like MC Dropout require warm-up epochs to yield meaningful uncertainty. Addressing this gap, post-hoc techniques such as that of Sourati et al. (2018), which enable uncertainty estimation from pretrained deterministic networks, represent a valuable and underexplored direction for early-stage sample selection.
7.5 Method suboptimality
To provide a necessary contrast to the benefits of many models, this section discusses key methodological limitations. These limitations will serve as guidance for readers selecting the most optimal methods and will be utilized when we present our recommendations in Section 8.
Invertible covariance of SSNs
Stochastic Segmentation Networks (SSNs), operate by placing a multivariate gaussian over the classification head of an existing model. However, they can suffer from training instability primarily due to the requirement that the covariance matrix remain invertible. To address this critical limitation, the authors propose two distinct remedial strategies: masking out the background to prevent exploding variances and employing uncorrelated gaussians when the covariance matrix becomes singular, which may impact model performance and theoretical purity.
GAN training
The Calibrated Adversarial Refinement network (CAR, as treated in Section 4.2.1), on the other hand, is a less commonly used method in literature. This can perhaps be accredited due to the well-known instability of GAN-based models (Arjovsky and Bottou, 2017). GANs often require additional heuristic terms in the loss function to ensure stable training. CAR is no exception to this, as its objective function combines four separate losses. The multiplicity of individual losses is critical to the model’s success. Nonetheless, this complex objective function suggests the training process may pose substantial fine-tuning challenges, potentially explaining the model’s limited practical adoption.
Mode collapse with ELBOs
Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, Variational Autoencoders (VAEs) seem to be the most popular choice for the task of encapsulating Observer Variability. (See Table 1 in Section 6.1). Nonetheless, the shortcomings of VAEs are well known. For example, such models suffer from inference suboptimality related to ELBO optimization (Cremer, 2018; Zhao et al., 2017) and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of mode collapse (Valiuddin et al., 2024a; b; Qiu and Lui, 2020). This has been hypothesized to be caused by excessively strong decoders (Chen et al., 2016) and is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective (Kohl et al., 2019a), residual connections (Kohl et al., 2019a; Baumgartner et al., 2019), or deep supervision (Baumgartner et al., 2019) are required for generalization.
Limitation of sequential modeling
The adoption of sequential sampling models DDPMs or PixelSeg is rather limited compared to the VAE-based models. This is unexpected, as both models outperforms the latter. This is likely due to their tedious sequential inference procedure (Song et al., 2020; Zheng et al., 2023; Meng et al., 2023; Zhang et al., 2022d), a crucial limitation that is exacerbated in supervised settings, which often require validation through sampling on a separate data split. Despite these shortcomings, the strengths of DDPMs should not be overlooked. Their iterative sampling allows for highly flexible modeling and preserves high-frequency details often lost in latent-variable approaches like VAEs, which can produce blurry reconstructions.
Discrete vs. continuous
It can be noted in Table 6 that the best-performing Denoising Diffusion Probabilistic Models (DDPMs) are discrete in nature (Chen et al., 2023a; Zbinden et al., 2023). While it is debatable whether shifting to categorical distributions is required for complex image generation (Chen et al., 2022). This can visually be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 14). However, Table 6 shows AB (Chen et al., 2022) performs almost identically to CCDM (Zbinden et al., 2023), suggesting that straightforward thresholding of continuous models have little-to-no performance sacrifice compared with explicitly modeling discrete distributions. These observations suggests that the modeling intuition of a categorical latent space, despite its fundamental soundness, lacks demonstrable merit.
(a) LIDC-IDRI (Armato, 2011).
(b) CityScapes (Cordts et al., 2016a).
Figure 14: Continuous vs. categorical forward diffusion process with cosine noise scheduling (Nichol and Dhariwal, 2021). Note the use of categorical diffusion that results in a more gradual transition for the multi-class case.
Criticism of MC Dropout
The majority of literature on parameter distribution modeling appears to rely on approximations of variational inference (VI). Among these, MC Dropout is the most commonly used due to its simplicity, low computational cost, and ease of implementation. However, MC Dropout has faced substantial criticism (Folgoc et al., 2021; Osband, 2016; Kingma et al., 2015; Gal et al., 2017a). For instance, it can assign zero probability to the true posterior or introduce erroneous multi-modality (Folgoc et al., 2021). It is also sensitive to model size and dropout rate rather than the observed data (Osband, 2016; Kingma et al., 2015). To address these issues, some alternatives attempt to learn the dropout rate (Kingma et al., 2015; Gal et al., 2017a). In cases of improved Model Generalization (Section 6.4), it is also debatable whether the performance gains stem from better uncertainty modeling or from other factors. For instance, both ensembling and MC Dropout introduce a model combination effect, effectively averaging multiple networks. MC Dropout also resembles placing an $L_{2}$ penalty on weights (Gal and Ghahramani, 2016). While these effects may improve generalization, similar gains might be achieved with simpler, more efficient regularizers. This suggests that perhaps it is the inherent stochasticity of the methods that is the key beneficial factor. Still, if achieving peak performance is the primary goal, uncertainty modeling is likely not the most efficient route due to its added computational cost. Instead, the improved performance may simply be an unintended side effect of techniques chosen for a different purpose.
7.6 Data dependency
In contrast to many other domains within computer vision, this field exhibits a notably greater absence of a single state-of-the-art method. Performance is heavily contingent upon multiple factors such as modeling context and application. However, data dependency often remains unexamined. Especially in cases where parameter distributions are modeled. For instance, MC dropout was found to perform better than ensembling (Hoebel et al., 2020; Roy et al., 2018b), while in other works ensembling excels (Mehrtash et al., 2020; Ng et al., 2022). Also, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout (Mukhoti and Gal, 2018). All things considered, the preference for a particular methodology seems to carry a strong data-dependency (Dahal et al., 2020; Jungo and Reyes, 2019; Burmeister et al., 2022; Ng et al., 2022; Kahl et al., 2024). For feature modeling, some models in Table 6 may appear to be overall best performers. However, their validation is often limited to binary, multi-annotated data, thereby compromising any claims of generalization to other data types, such as multi-class or single-annotated scenarios. Current research often focuses on incorporating uncertainty into already high-performing systems rather than comparing distinct methods. This trend necessitates a more rigorous, data-driven perspective to validate and understand the optimal uncertainty modeling strategy. Hence, a valuable benchmark study should comprehensively compare all available methods (including both feature and parameter modeling approaches) across all relevant task with special focus on the data dependencies.